Determine the speed of an object whose Kinect energy equals it's rest energy(2.6×10^8m/s)


1
Expert's answer
2021-07-14T10:15:36-0400

The Kinetic Energy of the object is

"K=\text{total energy - rest mass energy}\\\implies K= \dfrac{m_0c^2}{\sqrt{1-\frac{v^2}{c^2}}}-m_0c^2"


Since Kinetic Energy of the particle is equal to its rest energy. So,

"\implies \dfrac{m_0c^2}{\sqrt{1-\frac{v^2}{c^2}}}-m_0c^2=m_0c^2\\\ \\\implies \dfrac{m_0c^2}{\sqrt{1-\frac{v^2}{c^2}}}=2m_0c^2\\\ \\\implies \dfrac{1}{\sqrt{1-\frac{v^2}{c^2}}}=2\\\ \\\implies\dfrac{1}{4}=1-\dfrac{v^2}{c^2}\\\ \\\implies \dfrac{v^2}{c^2}=1-\dfrac{1}{4}=\dfrac{3}{4}\\\ \\\implies v=\dfrac{\sqrt 3}{2}c"


"\because c=2.6\times 10^8\ \ m/s"


So, speed of object (v) = "\dfrac{\sqrt 3}{2}\times 2.6\times 10^8 =2.25\times 10^8\ \ m/s"